# Fully non-linear elliptic equations on noncompact complex manifolds

### v1

@@ 3. Canonical metrics
Next, we calculate the length of the curve $\gamma(t)$
(3.8) $L(\gamma^{\prime}(t))$ $=\int_{0}^{1}\sqrt{\langle\gamma^{\prime}(t),\gamma^{\prime}(t)\rangle_{g}}{\rm d}t$
@@
(3.10) $L(\gamma^{\prime}(t))$ $\geq\int_{0}^{1}\frac{c}{\varphi}\frac{{\rm d}\varphi(\gamma(t))}{{\rm d}t}{\rm d}t$
@@
$M$ is a non-Kähler manifold, and $h_{i\bar{j}}$ is a Hermitian metric on $M$. Then there must exist a point $p\in M$ such that $h_{i\bar{j}}(p)$ is not a Kähler metric. Take a local holomorphic coordinate system chart $(V(p),(z^{1},\cdots,z^{n}))$ around $p$ and assume that $p$ is coordinate origin. Define $\varphi(z)=z\bar{z}-\epsilon$, where $\epsilon$ is a positive constant. Then if $\epsilon$ small enough, $\varphi$ can be seen as a defining function of a domain $\Omega:=\{z\in(V(p),(z^{1},\cdots,z^{n}))|\varphi(z)<0\}$ on $(M,h_{i\bar{j}})$, $\varphi$, $|\partial\varphi|_{h}$ and $|\partial\bar{\partial}\varphi|_{h}$ are bounded on $\Omega$, this is an example of Proposition 3.6.
@@ 4. $C^{0}$-estimates
Our main Theorem of this Section is as follows:
Theorem 4.1.
(4.1) $F(\alpha^{i\bar{p}}(\chi_{j\bar{p}}+u_{j\bar{p}}))=h(x,u)$
(4.3) $|u|\leq\zeta(1+\limsup_{z\rightarrow\infty}|u|).$
To prove this theorem, we first need the following lemma:
@@
To estimate $\sup_{\Omega}u$ from (4.10), we will give lower bound for $\mathcal{F}$, where $\mathcal{F}=\sum_{k}F^{kk}[\alpha^{i\bar{p}}(\chi+\sqrt{-1}u\partial\bar{\partial}\varphi)_{j\bar{p}}]$. We can choose a new coordinate system at $P$ such that $\alpha_{i\bar{j}}=\delta_{ij}$ and $(\chi+\sqrt{-1}u\partial\bar{\partial}\varphi)_{j\bar{p}}=a_{j}\delta_{jp}$, then $\alpha^{i\bar{p}}(\chi+\sqrt{-1}u\partial\bar{\partial}\varphi)_{j\bar{p}}]$ is a diagonal matrix, this follows that
@@
By (a5), we can choose $\sigma\in(\sup_{\Omega}h(x,0),\sup_{\Gamma}f)$, let $x$ be the closest point to the origin such that $f(x)=\sigma$. By the concavity of $f$ and symmetry under permuting the variables, we must have $x=N{\rm\textbf{1}}$ for some $N>0$. For any $\lambda\in\Gamma$, assumption (a4) implies that there is some $T>1$, such that $f(T\lambda)>\sigma$. By (a2) condition, we also have $f(Tx/(T-1))>\sigma$. The concavity of $f$ implies that then
@@ 5. $C^{1}$ and $C^{2}$-estimates
In the following, we consider the a priori estimates of the equations for $u\in C^{4}(M)\cap\hat{C}^{2}(M)$ as follows:
@@
Definition 5.1.
We say that $u$ is a $\mathcal{C}$-subsolution for the equation $F(A)=h$ if the following holds. Let $g_{i\bar{j}}$ be defined as in (5.2). We require that for every point $x\in M$, if $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denote the eigenvalues of the endomorphism $\alpha^{i\bar{p}}g_{j\bar{p}}$ at $x$, then for all $i=1,\ldots,n$ we have
(5.5) $\lim_{t\rightarrow\infty}f(\lambda+t\textbf{e}_{i})>h(x).$
@@
To use the generalized maximum principle, in the sequel, we assume that $u\in C^{4}(M)\cap\hat{C}^{2}(M)$. We consider the function
@@
The quantity $H$ need not be smooth at any point, because the largest eigenvalue of $A$ may have eigenspace of dimension larger than 1. Therefore, we can not apply Proposition 3.2 to $H$ directly. Since $u\in\tilde{C}^{4}(M)$, $H$ is bounded from above in $M$, there exists a a sequence $\{p_{i}\}$ in $M$ such that $\lim H(p_{i})=\sup H$. If $H$ achieves its maximum at some point $p\in M$, then we can use the same method as in [5] to obtain the estimate (5.7). Hence in the following, we assume that
@@
for a uniform $0<\beta<1$, where the norm is defined in (3.1) with respect to $(M,\alpha)$. Differentiating the equation and applying a standard bootstrapping argument, we finally obtain uniform higher-order estimates. Then we have the following theorem:
Theorem 5.2.
Let $(M,\alpha)$ be a complete Hermitian manifold with bounded geometry of order k1k-1, $k\geq 4$. Fix a real $(1,1)$-form $\chi$ with bounded geometry of order k1k-1 on $(M,\alpha)$, given $h\in\tilde{C}^{k-2+\beta}(M)$, let $u\in\tilde{C}^{k+\beta}(M)$ satisfy that
(5.50) $F(\alpha^{i\bar{p}}(\chi_{j\bar{p}}+u_{j\bar{p}}))=h,$
where $\beta\in(0,1)$ is some uniform constant. Suppose the $\underline{u}\in\tilde{C}^{\infty}(M)$ is a $\mathcal{C}$-subsolution for the equation (5.50). Suppose that the equation satisfies conditions (a1)-(a4). Then, we have an estimate $|u|_{k+\beta}\leq C_{k}$, with constant $C_{k}$ depending on $k$, on the background date $M,\alpha,\chi,F,h$, on $\sup_{M}|u|$, and the subsolution $\underline{u}$.
@@ 6. The solvability of $F(A)=h$
and since $\sqrt{-1}\partial\bar{\partial}\varphi>0$, by (6.5), (6.6) and the fact the $F$ is concave, there exists a constant $c_{1}>0$ such that
(6.7) $\frac{1}{c_{1}}\delta_{ij}\varphi_{i\bar{j}}\leq\sum A^{i\bar{j}}\varphi_{i\bar{j}}\leq c_{1}\delta_{ij}\varphi_{i\bar{j}}.$
@@
Next, we set up the continuity method to solve the equation (6.10) as in Theorem 6.2, we assume that
@@
Closedness. By Theorem 6.2, we obtain the closedness of $\mathcal{L}$.
@@
Theorem 6.3.
Assume that $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, then the equation (6.10) as in Theorem 6.2 exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$.
Finally, we obtain the following existence result for the solution of $F(A)=h$.
Theorem 6.4.
Let $\Omega$ be a complete Kähler manifold with a bounded global Kähler potential $\varphi<0$ and $\lim_{x\rightarrow\infty}\varphi(x)=0$. Let $\alpha$ be a Hermitian metric on $\Omega$ with bounded geometry of order k1k-1, where $k\geq 4$. Fix a real $(1,1)$-form $\chi$ with bounded geometry of order k1k-1, and $\chi$ is uniformly equivalent to $\alpha$ on $(\Omega,\alpha)$, given $h\in\tilde{C}^{k-2+\beta}(\Omega\times\mathbb{R})$, where $\beta\in(0,1)$ is some uniform constant. Consider the following equation on $\Omega$:
(6.30) $F(\alpha^{i\bar{p}}(\chi_{j\bar{p}}+u_{j\bar{p}}))=h(x,u),$
Suppose that the above equation satisfies conditions (a1)-(a5) and $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, assume that $|h(x,0)|=O(\underset{\min}{\lambda}(\alpha^{a\bar{l}}\varphi_{b\bar{l}}))$, and for all $i=1,\ldots,n$
$\textbf{\emph{e}}_{i}$ denotes the $i$th standard basis vector. Then the equation (6.30) exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$ such that $\chi+\sqrt{-1}\partial\bar{\partial}u$ is uniformly equivalent to $\alpha$.
Proof.
@@
By Theorem 6.2 and Theorem 6.3, for each $m$, the equation (6.32) exists a solution $u_{m}\in\tilde{C}^{k+\beta}(\Omega)$ and $|u_{m}|_{k+\beta}<c$ for some constant $c$ independent of $m$. Let $\{\Omega_{i}\}$ be an exhaustion of $\Omega$ by compact subdomains. By the Arzelà-Ascoli Theorem, on $\Omega_{1}$ we can take a subsequence $\{u_{m_{k}}\}$ of $\{u_{m}\}$ such that $u_{m_{k}}\rightarrow u^{1}$ for some $u^{1}\in\tilde{C}^{k+\beta}(\Omega_{1})$. And on $\Omega_{2}$, we can also take a subsequence of $\{u_{m_{k}}\}$ converging to some $u^{2}\in\tilde{C}^{k+\beta}(\Omega_{2})$. Note that $u^{2}|_{\Omega_{1}}=u^{1}$. Continuing in this way, we obtain a function sequence $u^{i}\in\tilde{C}^{k+\beta}(\Omega_{i})$ such that $u^{i}|_{\Omega_{i-1}}=u^{i-1}$ and $|u_{i}|_{k+\beta}<c$ for some constant $c$ independent of $i$. Define $u\in\tilde{C}^{k+\beta}(\Omega)$ by $u|_{\Omega_{i}}=u^{i}$ for each $i$, we can verify that $u$ satisfies equation (6.30). ∎
@@
Theorem 6.5.
Let $(\Omega,\alpha)$ be a complete Hermitian manifold with bounded geometry of order k1k-1, where $k\geq 4$. Fix a real $(1,1)$-form $\chi$ with bounded geometry of order k1k-1, and $\chi$ is uniformly equivalent to $\alpha$ on $(\Omega,\alpha)$, given $h\in\tilde{C}^{k-2+\beta}(\Omega\times\mathbb{R})$ such that $h_{u}\geq a>0$ for some constant $a$, where $\beta\in(0,1)$ is some uniform constant. Consider the following equation on $\Omega$:
(6.33) $F(\alpha^{i\bar{p}}(\chi_{j\bar{p}}+u_{j\bar{p}}))=h(x,u),$
Suppose that the above equation satisfies conditions (a1)-(a5) and $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, and for all $i=1,\ldots,n$
$\textbf{\emph{e}}_{i}$ denotes the $i$th standard basis vector. Then the equation (6.33) exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$ such that $\chi+\sqrt{-1}\partial\bar{\partial}u$ is uniformly equivalent to $\alpha$.
Proof.
The proof is similar to that of Theorem 6.3. We only need to use Theorem 4.4 instead of Theorem 4.1 to obtain the $C^{0}$ estimates for (6.33), the rest of the proof is the same as that of Theorem 6.3. ∎
7. Monge-Ampère equation
@@
the equation (1.1) is the Monge-Ampère equation. Let (6.30) be a Monge-Ampère equation, then by Theorem 6.4, we obtain the following:
@@
Let $\Omega$ be a strictly pseudoconvex domain in a complete Kähler manifold $(M,h_{i\bar{j}})$ as in Proposition 3.5 with $l\geq 3$. Define $g=-\log(-\varphi)$ and $\alpha_{i\bar{j}}=g_{i\bar{j}}+\hat{h}_{i\bar{j}}$, where $\varphi$ is the defining function of $\Omega$. Given $h\in\tilde{C}^{l-1+\beta}(\Omega)$, where $\beta\in(0,1)$ is some uniform constant. Consider the following equation on $\Omega$:
@@
Next, we will consider the (n1)(n-1) Monge-Ampère equation. We introduce some basic properties about (n1,n1)(n-1,n-1) forms:
(7.7) $\omega\mapsto\omega^{n-1}.$
(7.8) $\Theta=$ $(\sqrt{-1})^{n-1}(n-1)!$
@@
(7.9) $s(p,q)=\left\{\begin{aligned} -1,\;\;\;\;{\rm if}\;p>q;\\ 1,\;\;\;\;\;{\rm if}\;p\leq q.\end{aligned}\right$.
Here we introduce the sign function $s$ such that,
(7.10) $dz^{p}$ $\wedge d\bar{z}^{q}\wedge s(p,q)\wedge dz^{1}\wedge d\bar{z}^{1}\cdots\wedge\widehat{dz^{p}}\wedge d\bar{z}^{p}\wedge\cdots\wedge dz^{q}\wedge\widehat{d\bar{z}^{q}}\cdots\wedge dz^{n}\wedge d\bar{z}^{n}$
We denote
(7.11) $\det\Theta=\det(\Theta_{p\bar{q}}).$
$\sum_{l}\Theta_{i\bar{l}}\Theta^{j\bar{l}}=\delta_{ij}.$
We note that, for a positive $(1,1)$-form $\omega$ given by
(7.12) $\omega=\sqrt{-1}\sum_{i,j=1}^{n}g_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j},$
we obtain
(7.13) $\omega^{n}=(\sqrt{-1})^{n}n!\det(g_{i\bar{j}})dz^{1}\wedge d\bar{z}^{1}\cdots\wedge dz^{n}\wedge d\bar{z}^{n},$
and we have
(7.14) $(\omega^{n-1})_{i\bar{j}}=\det(g_{i\bar{j}})g^{i\bar{j}}.$
It follows that
(7.15) $\det(\omega^{n-1})=\det(g_{i\bar{j}})^{n-1}.$
Then, we consider the form-type (n1)(n-1)-Monge-Ampère equation on $M$ as follows:
@@
where $\omega$ is a Kähler form as in (7.12) and $\omega_{0}$ is the form of a balanced metric on $M$.
Write * for the Hodge star operator with respect to $\omega$. The Hodge star operator * takes any (n1,n1)(n-1,n-1) form to a $(1,1)$ form, and vice versa. This acts on real (n1,n1)(n-1,n-1)-forms as follows. Consider a real (n1,n1)(n-1,n-1) given by (7.8), if we are computing at a point in coordinates so that $\omega_{i\bar{j}}=\delta_{ij}$, then
(7.17) $\Theta=\sqrt{-1}(n-1)!\sum_{p,q}\Theta_{p\bar{q}}dz^{q}\wedge d\bar{z}^{p},$
and
(7.18) $\Theta=\Theta.$
Then by (7.13)-(7.18), we have
@@
and
@@
Let $(M,\alpha)$ be a Hermitian manifold of complex dimension $n$ and write
(7.24) $\alpha=\sqrt{-1}\alpha_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}>0.$
(7.25) $g_{i\bar{j}}:=\chi_{i\bar{j}}+u_{i\bar{j}}.$
Let us write
(7.26) $\tilde{g}_{i\bar{j}}:=P_{\alpha}(g_{i\bar{j}})=\frac{1}{n-1}(({\rm tr}_{\alpha}g)\alpha_{i\bar{j}}-g_{i\bar{j}}),$
where $P_{\alpha}$ is an operator on tensors, depending on the fixed metric $\alpha$, defined by the second equality in (7.26). Here ${\rm tr}_{\alpha}g$ is the trace of $g$ with respect to $\alpha$. As an aside, if $\alpha$ is the Euclidean metric on $\mathbb{C}^{n}$, then the condition $P_{\alpha}(u_{i\bar{j}})\geq 0$ is equivalent to saying that $u$ is (n1)(n-1)-plurisubharmonic. Observe that, writing $\Delta=\alpha^{k\bar{l}}\partial_{k}\partial_{\bar{l}}$,
(7.27) $\tilde{g}_{i\bar{j}}=\tilde{\chi}_{i\bar{j}}+\frac{1}{n-1}((\Delta u)\alpha_{i\bar{j}}-u_{i\bar{j}}).$
Given a smooth function $h$, we consider equations of the form
(7.28) $F(A)=h,$
where $A$ is the endomorphism $A_{j}^{i}=\alpha^{i\bar{p}}g_{j\bar{p}}$ of the holomorphic tangent bundle, which is Hermitian with respect to the inner product defined by $\alpha$, and $F(A)$ is a symmetric function of the eigenvalues $\lambda_{1},\ldots,\lambda_{n}$ of $A$:
(7.29) $F(A)=f(\lambda_{1},\ldots,\lambda_{n}).$
We assume that our operator $F$ has the special form $F(M)=\widetilde{F}(P(M))$, where
@@
analogous to $P_{\alpha}$ above (here writing ${\rm Tr}(M)$ for the trace of matrix $M$), and
(7.31) $\widetilde{F}(B)=\tilde{f}(\mu_{1},\ldots,\mu_{n}),$
where $\mu_{1},\ldots,\mu_{n}$ are the eigenvalues of $B$, and $\tilde{f}$ is another symmetric function. In terms of eigenvalues, this means that
(7.32) $f(\lambda_{1},\ldots,\lambda_{n})=(\tilde{f}\circ P)(\lambda_{1},\ldots,\lambda_{n}),$

### v2

@@ 3. Canonical metrics
Next, we calculate the length of the curve $\gamma(t)$, which we denote by $L(\gamma(t))$,
(3.7) $L(\gamma(t))$ $=\int_{0}^{1}\sqrt{\langle\gamma^{\prime}(t),\gamma^{\prime}(t)\rangle_{g}}{\rm d}t$
@@
(3.9) $L(\gamma(t))$ $\geq\int_{0}^{1}\frac{c}{\varphi}\frac{{\rm d}\varphi(\gamma(t))}{{\rm d}t}{\rm d}t$
@@
$M$ is a non-Kähler manifold, and $h_{i\bar{j}}$ is a Hermitian metric on $M$. Then there must exist a point $p\in M$ such that $h_{i\bar{j}}(p)$ is not a Kähler metric. Take a local holomorphic coordinate system chart $(V(p),(z^{1},\cdots,z^{n}))$ around $p$ and assume that $p$ is coordinate origin. Define $\varphi(z)=z\bar{z}-\epsilon$, where $\epsilon$ is a positive constant. Then if $\epsilon$ small enough, $\varphi$ can be seen as a defining function of a domain $\Omega:=\{z\in(V(p),(z^{1},\cdots,z^{n}))|\varphi(z)<0\}$ on $(M,h_{i\bar{j}})$, $\varphi$, $|\partial\varphi|_{h}$ and $|\partial\bar{\partial}\varphi|_{h}$ are bounded on $\Omega$.
@@ 4. $C^{0}$-estimates
Our main Theorem of this Section is Theorem 1.2. To prove this theorem, we first need the following lemma:
@@
To estimate $\sup_{\Omega}u$ from (4.7), we will give lower bound for $\mathcal{F}$, the method is similar to Lemma 9 in [25], where $\mathcal{F}=\sum_{k}F^{kk}[\alpha^{i\bar{p}}(\chi+\sqrt{-1}u\partial\bar{\partial}\varphi)_{j\bar{p}}]$. We can choose a new coordinate system at $P$ such that $\alpha_{i\bar{j}}=\delta_{ij}$ and $(\chi+\sqrt{-1}u\partial\bar{\partial}\varphi)_{j\bar{p}}=a_{j}\delta_{jp}$, then $\alpha^{i\bar{p}}(\chi+\sqrt{-1}u\partial\bar{\partial}\varphi)_{j\bar{p}}]$ is a diagonal matrix, this follows that
@@
By (a5), we can choose $\sigma\in(\sup_{\Omega}h(x,0),\sup_{\Gamma}f)$, let $x$ be the closest point to the origin such that $f(x)=\sigma$. By the concavity of $f$ and symmetry under permuting the variables, we must have $x=N{\rm\textbf{1}}$ for some $N>0$, where ${\rm\textbf{1}}=(1,\ldots,1)\in\mathbb{R}^{n}$. For any $\lambda\in\Gamma$, assumption (a4) implies that there is some $T>1$, such that $f(T\lambda)>\sigma$. By (a2) condition, we also have $f(Tx/(T-1))>\sigma$. The concavity of $f$ implies that then
@@
In the following, we consider the a priori estimates of the equations for $u\in C^{4}(M)\cap\{u\in C^{2}(M)|\;|u|_{2}<\infty\}$ as follows:
@@
To use the generalized maximum principle, in the sequel, we assume that $u\in C^{4}(M)\cap\{u\in C^{2}(M)|\;|u|_{2}<\infty\}$. We consider the function
@@
The quantity $H$ need not be smooth at any point, because the largest eigenvalue of $A$ may have eigenspace of dimension larger than 1. Therefore, we can not apply Proposition 3.2 to $H$ directly. Since $u\in C^{4}(M)\cap\{u\in C^{2}(M)|\;|u|_{2}<\infty\}$, $H$ is bounded from above in $M$, there exists a a sequence $\{p_{i}\}$ in $M$ such that $\lim H(p_{i})=\sup H$. If $H$ achieves its maximum at some point $p\in M$, then we can use the same method as in [24] to obtain the estimate (5.6). Hence in the following, we assume that
@@
for a uniform $0<\beta<1$, where the norm is defined in (3.1) with respect to $(M,\alpha)$. Differentiating the equation and applying a standard bootstrapping argument, we finally obtain uniform higher-order estimates. Then we can prove Theorem 1.4.
@@ 6. The solvability of $F(A)=h$
this means that
(6.7) $\{\lambda(\alpha^{i\bar{p}}(\chi_{j\bar{p}}+tu_{j\bar{p}})):t\in[0,1]\}\subset K\subset\Gamma_{n},$
where $K$ is some bounded closed set in $\mathbb{R}^{n}$, and since $\sqrt{-1}\partial\bar{\partial}\varphi>0$, by (6.5), (6.7) and the condition (a1), there exists a constant $c_{1}>0$ such that
(6.8) $\frac{1}{c_{1}}\sum\delta_{ij}\varphi_{i\bar{j}}\leq\sum A^{i\bar{j}}\varphi_{i\bar{j}}\leq c_{1}\sum\delta_{ij}\varphi_{i\bar{j}}.$
@@
Next, we set up the continuity method to solve the equation (6.11) as in Lemma 6.2, we assume that
@@
Closedness. By Lemma 6.2, we obtain the closedness of $\mathcal{L}$.
@@
Lemma 6.3.
Assume that $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, then the equation (6.11) as in Lemma 6.2 exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$.
Finally, we obtain the following existence result (Theorem 1.5) for the solution of $F(A)=h$.
Proof of Theorem 1.5.
@@
By Lemma 6.2 and Lemma 6.3, for each $m$, the equation (6.31) exists a solution $u_{m}\in\tilde{C}^{k+\beta}(\Omega)$ and $|u_{m}|_{k+\beta}<c$ for some constant $c$ independent of $m$. Let $\{\Omega_{i}\}$ be an exhaustion of $\Omega$ by compact subdomains. By the Arzelà-Ascoli Theorem, on $\Omega_{1}$ we can take a subsequence $\{u_{m_{k}}\}$ of $\{u_{m}\}$ such that $u_{m_{k}}\rightarrow u^{1}$ for some $u^{1}\in\tilde{C}^{k+\beta}(\Omega_{1})$. And on $\Omega_{2}$, we can also take a subsequence of $\{u_{m_{k}}\}$ converging to some $u^{2}\in\tilde{C}^{k+\beta}(\Omega_{2})$. Note that $u^{2}|_{\Omega_{1}}=u^{1}$. Continuing in this way, we obtain a function sequence $u^{i}\in\tilde{C}^{k+\beta}(\Omega_{i})$ such that $u^{i}|_{\Omega_{i-1}}=u^{i-1}$ and $|u_{i}|_{k+\beta}<c$ for some constant $c$ independent of $i$. Define $u\in\tilde{C}^{k+\beta}(\Omega)$ by $u|_{\Omega_{i}}=u^{i}$ for each $i$, we can verify that $u$ satisfies equation (1.10). ∎
@@
Proof of Theorem 1.6.
The proof is similar to that of Lemma 6.3. We only need to use Theorem 4.3 instead of Theorem 1.2 to obtain the $C^{0}$ estimates for (1.11), the rest of the proof is the same as that of Lemma 6.3. ∎
7. Complex Monge-Ampère equations
@@
where $f$ is defined on $\Gamma_{n}$, the equation (1.1) is the Monge-Ampère equation. It is straightforward to check that $f$ satisfies the structure conditions for equation (1.1). (a1), (a2) and (a5) are obvious. Indeed, $f$ converges to -$\infty$ on the boundary $\partial\Gamma_{n}$, so (a3) is satisfied, and for (a4) it is enough to note that $f(t\mu)=f(\lambda)+n\log t$, which converges to $\infty$ as $t\rightarrow\infty$.
In addition, if $\lambda\in\Gamma_{n}$, then we also have
(7.2) $\lim_{t\rightarrow\infty}f(\lambda+t\textbf{e}_{i})=\infty$
for all $i$. This means that for a function $u$ to be a $\mathcal{C}$-subsolution for this equation, the only requirement is that $\lambda[\alpha^{i\bar{p}}\chi_{j\bar{p}}]\in\Gamma_{n}$ (see Page 1).
Let (1.10) be a Monge-Ampère equation, then by Theorem 1.5, we obtain the following:
@@
Let $\Omega$ be a strictly pseudoconvex domain in a complete Kähler manifold $(M,\hat{h}_{i\bar{j}})$ as in Proposition 3.5 with $l\geq 3$. Define $g=-\log(-\varphi)$ and $\alpha_{i\bar{j}}=g_{i\bar{j}}+\hat{h}_{i\bar{j}}$, where $\varphi$ is the defining function of $\Omega$. Given $h\in\tilde{C}^{l-1+\beta}(\Omega)$, where $\beta\in(0,1)$ is some uniform constant. Consider the following equation on $\Omega$:
@@
The above corollary can be used to find Kähler metrics with some prescribed volume forms on strictly pseudoconvex domains, and we can prove Theorem 1.7.
Proof of Theorem 1.7.
Since $|\partial\bar{\partial}\varphi|_{h}>c>0$, we obtain that $(\Omega,g_{i\bar{j}})$ has bounded geometry of infinity order. Note that
(7.4) $\varphi^{2(1-n)}=O(\varphi^{2})|(\sqrt{-1})^{n}\det(g_{i\bar{j}})dz^{1}\wedge d\bar{z}^{1}\cdots\wedge dz^{n}\wedge d\bar{z}^{n}|_{h},$
then by (1.12), we have
(7.5) $\frac{|V(x)-\det(g_{i\bar{j}})|}{\det(g_{i\bar{j}})}=O(\varphi^{2}).$
Finally,
(7.6) $V(x)=e^{\tilde{h}}\det(g_{i\bar{j}}),$
where $\tilde{h}\in\tilde{C}^{\infty}(\Omega)$ and $|\tilde{h}|=O(\varphi^{2})$. Consider the following equation
(7.7) $\det(g_{i\bar{j}}+u_{i\bar{j}})=V(x)=e^{\tilde{h}}\det(g_{i\bar{j}}),$
taking $g_{i\bar{j}}$ instead of $\alpha_{i\bar{j}}$ in (7.3) of Corollary 7.1, note that $\underset{\min}{\lambda}(h^{a\bar{j}}\varphi_{b\bar{j}})>c^{\prime}>0$, we obtain that (7.7) has a solution $u\in\tilde{C}^{\infty}(\Omega)$ such that $(g_{i\bar{j}}+u_{i\bar{j}})$ is uniformly equivalent to $\alpha$. This completes the proof of the theorem. ∎
@@
Indeed, Corollary 7.2 can be used to find Kähler-Einstein metrics on some submanifolds of a complete Kähler manifold whose first Chern class is negative. The following functions $f$, $\psi$ and $\mathfrak{F}$ can be found in [20]:
Let $\kappa\in(0,1)$, $f:[0,1)\rightarrow[0,\infty)$ be the function:
(7.12) $f(s)=\left\{\begin{aligned} &0,\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;s\in[0,1-\kappa];\\ &-\log\biggl[1-\biggl(\frac{s-1+\kappa}{\kappa}\biggl)^{2}\biggl],s\in(1-\kappa,1).\end{aligned}\right$.
Let $\psi\geq 0$ be a smooth function on $\mathbb{R}^{+}$ such that
(7.13) $\psi(s)=\left\{\begin{aligned} &0,\;\;\;s\in[0,1-\kappa+\kappa^{2}];\\ &1,\;\;\;s\in(1-\kappa+2\kappa^{2},1).\end{aligned}\right$.
and $\frac{2}{\kappa^{2}}\geq\psi^{\prime}\geq 0$. Define
$\mathfrak{F}:=\int_{0}^{s}\psi(\tau)f^{\prime}(\tau)d\tau$
By [20], we have
Lemma 7.3.
Suppose $0<\kappa<\frac{1}{8}$. Then the function $\mathfrak{F}\geq 0$ defined above is smooth and satisfies the following:
   (i) $\mathfrak{F}=0$ for $s\in[0,1-\kappa+\kappa^{2}]$.
  (ii) $\mathfrak{F}^{\prime}\geq 0$ and for any $k\geq 1$, $\exp(-k\mathfrak{F})\mathfrak{F}^{(k)}$ is uniformly bounded.
Then, we can prove Theorem 1.8, we rewrite it here:
Theorem 7.4.
Let $(M,\alpha)$ be a complete Kähler manifold with negative Ricci curvature bounded from below by a negative constant. Then, for any compact subset $K\subset M$, there exists a Kähler-Einstein metric on $K$. Moreover, this metric can be extended to some open set $N\subset M$ and forms a complete Hermitian metric on $N$.
Proof.
Since that ${\rm Ric}(\alpha)$ is bounded from below by a negative constant, by Shi [26] (see also Lemma 3.2 of [17]), there exists a smooth real function $\rho$ which is uniformly equivalent to the distance function from a fixed point. For any $\rho_{0}>0$, let $U_{\rho_{0}}$ be the component of $\{x|\rho(x)<\rho_{0}\}$ that contains the fixed point as above. Hence $U_{\rho_{0}}$ will exhaust $M$ as $\rho_{0}\rightarrow\infty$. For $\rho_{0}>>1$, we define $F(x)=\mathfrak{F}(\rho(x)/\rho_{0})$. Let $h_{0}=-e^{2F}{\rm Ric}(\alpha)$. We find $(U_{\rho_{0}},h_{0})$ is a complete Hermitian metric with bounded geometry of infinite order (see [20]) and $h_{0}=-{\rm Ric}(\alpha)$ on $\{\rho(x)<(1-\kappa+\kappa^{2})\rho_{0}\}$. If we choose $\rho_{0}$ large enough, we can assume the $K\subset\{\rho(x)<(1-\kappa+\kappa^{2})\rho_{0}\}$. Consider the following equation on $U_{\rho_{0}}$:
(7.14) $\det((h_{0})_{i\bar{j}}+u_{i\bar{j}})=e^{u}e^{h}\det((h_{0})_{i\bar{j}}),$
@@
(7.15) $h:=\log\frac{\det\alpha}{\det(-{\rm Ric}(\alpha))}.$
Since $\overline{U_{\rho_{0}}}$ is compact, the Kähler metric ${\rm Ric}(\alpha)$ has bounded geometry of infinity order on $U_{\rho_{0}}$, and $h\in\tilde{C}^{\infty}(U_{\rho_{0}})$ with respect to ${\rm Ric}(\alpha)$. By Lemma 7.3, we have $h_{0}\geq-{\rm Ric}(\alpha)$, hence we also have $h\in\tilde{C}^{\infty}(U_{\rho_{0}})$ with respect to $h_{0}$. Then by Corollary 7.2, the equation (7.14) exists a solution $u\in\tilde{C}^{\infty}(U_{\rho_{0}})$ such that $((h_{0})_{i\bar{j}}+u_{i\bar{j}})$ is uniformly equivalent to $((h_{0})_{i\bar{j}})$. By (7.14) and (7.15), we obtain
(7.16) ${\rm Ric}((h_{0})_{p\bar{q}}+u_{p\bar{q}})_{i\bar{j}}$ $=-\frac{\partial^{2}}{\partial z^{i}\partial\bar{z}^{j}}[\log\det((h_{0})_{p\bar{q}}+u_{p\bar{q}})]$
$=-\frac{\partial^{2}}{\partial z^{i}\partial\bar{z}^{j}}[u+\log\frac{\det\alpha}{\det(-{\rm Ric}(\alpha))}+\log\det((h_{0})_{p\bar{q}})]$
$=-((h_{0})_{i\bar{j}}+u_{i\bar{j}}),$
on $\{\rho(x)<(1-\kappa+\kappa^{2})\rho_{0}\}$. Finally, $((h_{0})_{i\bar{j}}+u_{i\bar{j}})$ is a Kähler-Einstein metric on $K$ and it is a complete Hermitian metric on $N:=U_{\rho_{0}}$. ∎
Next, we consider the form-type (n1)(n-1)-Monge-Ampère equation on $M$, which can be seen as another example of (1.1). Some basic properties about (n1,n1)(n-1,n-1) forms can be found in Appendix A.
@@
where $\omega$ is a Kähler form as in (A.6) and $\omega_{0}$ is the form of a balanced metric on $M$.
Then by (A.7)-(A.11), we have
@@
where * is the Hodge star operator with respect to $\omega$, and
@@
We assume that our operator $F$ (see (1.1)) has the special form $F(M)=\widetilde{F}(P(M))$, where
@@
(here writing ${\rm Tr}(M)$ for the trace of matrix $M$), and
(7.24) $\widetilde{F}(B)=\log(\mu_{1},\ldots,\mu_{n}),$
where $\mu_{1},\ldots,\mu_{n}$ are the eigenvalues of $B$. In terms of eigenvalues, this means that
(7.25) $f(\lambda_{1},\ldots,\lambda_{n})=(\log\circ P)(\lambda_{1},\ldots,\lambda_{n}),$